# Dark-EvGS: Event Camera as an Eye for Radiance Field in the Dark

### v1

@@ Abstract
In low-light environments, conventional cameras often struggle to capture clear multi-view images of objects due to dynamic range limitations and motion blur caused by long exposure. Event cameras, with their high-dynamic range and high-speed properties, have the potential to mitigate these issues. Additionally, 3D Gaussian Splatting (GS) enables radiance field reconstruction, facilitating bright frame synthesis from multiple viewpoints in low-light conditions. However, naively using an event-assisted 3D GS approach still faced challenges because, in low light, events are noisy, frames lack quality, and the color tone may be inconsistent. To address these issues, we propose Dark-EvGS, the first event-assisted 3D GS framework that enables the reconstruction of bright frames from arbitrary viewpoints along the camera trajectory. Triplet-level supervision is proposed to gain holistic knowledge, granular details, and sharp scene rendering. The color tone matching block is proposed to guarantee the color consistency of the rendered frames. Furthermore, we introduce the first real-captured dataset for the event-guided bright frame synthesis task via 3D GS-based radiance field reconstruction. Experiments demonstrate that our method achieves better results than existing methods, conquering radiance field reconstruction under challenging low-light conditions. The code and sample data are included in the supplementary material.
1 Introduction
@@
Low-light radiance field reconstruction plays a crucial role in various real-world applications, including nighttime photography [16], surveillance [15], and autonomous driving [8]. However, traditional frame-based cameras struggle to capture high-quality images in dark environments due to their limited dynamic range and reliance on long exposure times. These constraints result in significant noise and motion blur, making it difficult to capture clear, bright frames of objects from multiple viewpoints in low-light conditions.
Event cameras are a novel type of sensor that captures brightness changes asynchronously per pixel, marking a shift from conventional frame-based imaging [38, 29, 37]. With high dynamic range (HDR) and high temporal resolution capabilities, they can simultaneously capture bright and dark regions with no motion blur, producing rich unseen signals omitted by frame-based cameras, providing a potential solution for reconstructing in low-light conditions. Event-based video reconstruction methods were proposed [20, 6, 14] to output consecutive frames from events. Nevertheless, either they are limited to grayscale images [20, 6], or primarily focus on the short exposure problem [14] rather than low-light conditions (less than 40 lux). Event data can also be applied to radiance field reconstruction, but existing approaches [25, 21, 27] are only designed for normal light conditions (around 300 lux).
Reconstructing a radiance field in the dark is a hard task because of three main challenges: 1) Signals captured lack quality as events are noisier and frames contain lower intensity (Fig. 1 (a) and (c)); 2) color tone consistency is required for the rendered frames; and 3) there are currently no event-based dark radiance field reconstruction work and dataset available so no prior solution and data can be referenced for solving this problem. Due to these challenges, not only did the existing event-based radiance field approach [25, 21, 27] fail to become an effective solution, but also, event-based video reconstruction work [20, 6, 14] cannot be used directly in this task.
In this paper, we propose Dark-EvGS by integrating the event camera and 3D Gaussian Splatting (GS) [12], a radiance field reconstruction technique, enabling bright frame reconstruction from arbitrary viewpoints along the camera trajectory. The high-dynamic sensing capability of event cameras enables Dark-EvGS to have precise perception of camera motion, which can provide valuable guidance for radiance field reconstruction performance (Fig. 1 (b)) and render bright frames from multiple viewpoints of objects in low-light conditions (Fig. 1 (d)).
@@
  • We introduce Dark-EvGS, the first event-guided bright frame synthesis method via 3D GS-based radiance field reconstruction from arbitrary viewpoints in dark environments.
@@
  • We collect the first real-world event-based low-light radiance field reconstruction dataset with paired low-light, bright-light frames, event streams, and corresponding camera poses. We will release the dataset and code to facilitate future research.
2 Related Work
@@
Most NeRF and 3D GS-based methods use frame data from traditional cameras [18]. Few works have been attempted to reconstruct radiance fields in dark and low-light environments due to the limited dynamic range of a frame-based camera. Zhang et al. [35] attempt to solve robot exploration in the dark using 3D Gaussians Relighting, however, a constant illumination is placed in front of the robot. Mildenhall et al. [18] use NeRF to reconstruct the radiance field in the dark by treating it as a denoising problem from HDR frames. However, it requires an HDR sensor in real applications and does not explore the effect of motion blur in low-light conditions. Therefore, no existing work has attempted to leverage the dynamic range and temporal resolution property of an event camera to solve such an issue.
Event-Based Image Reconstruction. Many works have been devoted in the field of events to frame reconstruction [20, 1, 6]. However, these methods do not work well for reconstructing frames from low-light events. EvLowLight [14] attempts to reconstruct bright-light video by incorporating event data. However, the focus is primarily on the short-to-long exposure problem rather than the dark-to-bright reconstruction. In their case, the darkness of the frames is caused by short exposure times, not actual low-light conditions, meaning the scene remains well-lit from the event camera’s perspective. In contrast, our task focuses on using an event camera to reconstruct bright radiance fields in real low-light environments. Others [34, 19] have tried to use an event to guide video reconstruction in the dark. However, estimating absolute intensity values in a video solely from brightness changes recorded in events is a highly ill-posed problem [14]. Thus, directly applying any of these models to our task is not feasible.
@@
3 Methodology
3.1 Preliminary on 3D Gaussian Splatting
Figure 2: Overview of the Dark-EvGS pipeline for radiance field reconstruction in the dark. We obtain dark frames, events, and camera parameters using an event camera under low lights. The frame encoder and event encoder will extract features, which will be then forwarded to multimodal coherence modeling. The proposed Color Tone Consistency Module takes the features and decodes them into pseudo bright frames with consistent color tone (described in Sec. 3.3). Via the 3D GS model (described in Sec. 2), we render bright and shape frames when given the corresponding camera position and pose and formulate (blue arrow) the rendered results to predicted event signals (described in Sec. 3.2). The same process applies when formulating captured events to ground truth supervision signals. To supervise accurate training, the proposed triplet-level loss (red lines) provides a holistic view while keeping the rendered results sharp and accurate in detail (described in Sec. 3.4). As a result, our method enables high-quality radiance field reconstruction in the dark.
@@ 3.1 Preliminary on 3D Gaussian Splatting
3.2 Noisy Events in Dark as Supervisory Signals
Figure 3: Frame-based cameras are unable to capture enough signals for residence reconstruction in the dark. Event cameras, on the other hand, can capture ample signals, including ignored details by frame-based cameras (a). However, these raw events captured are rather noisy and random (a), which is one of the biggest challenges for existing methods [14, 25, 21, 20]. To solve this problem, we proposed a utilization (Sec. 3.2) for noisy events for accurate and clean supervisory signals (b).
@@ 3.2 Noisy Events in Dark as Supervisory Signals
Events triggered in low-light environments are much noisier [11]. To address noise, which is often more prevalent in microsecond-level event data [28], we preprocess the event stream using the $y$-noise filter from [7]. As shown in Fig. 3 (a), raw events captured in the dark contain extreme noise and randomness. After applying the noise filter, the cleaned event stream becomes more helpful as a supervision signal, as shown in Fig. 3 (b). Noise filtering is essential for accurate view rendering, as shown by ablation studies in the supplementary material.
Our objective, depicted in Fig. 2, is to reconstruct a radiance field using differentiable 3D Gaussian functions $G$, guided by event and corresponding blurred dark frames. We accumulate the ground truth event signal between timestamps $t_{1}$ and $t_{2}$ to form the supervisory signal using Eq. 5 by aggregating the polarities of all events occurring between times $t$ and $t-w$, indexed by their pixel location.
@@
3.3 Color Tone Consistency Module
Pure event utilization and supervision are insufficient for radiance field reconstruction on real data [9, 25], let alone reconstruction under low-light environments. Therefore, prior knowledge is required in the 3D GS training pipeline. Following EvLowLight [14], we deploy feature encoders for both frame and event modality along with the Multimodal Coherence Modeling block (MCM). We then propose the Color Tone Consistency Module (CTCM), which consists of a proposed Color Tone Matching Block (CTMB) along with the decoders from MCM. CTMB takes the decoded frames during the decoding process and corrects the color tone. Finally, the CTCM outputs the predicted pseudo frames using the corrected frames and the exposure parameter predicted from the MCM. Specifically, the CTMB, inspired by recent advancements in transposed self-attention [24], takes the last feature map $F\in\mathbb{R}^{H\times W\times C}$, and derives query $Q$, key $K$, and value $V$ representations using a 1×1 convolution followed by a depth-wise convolution. Subsequently, $Q$ and $V$ are reshaped, and an attention map $M\in\mathbb{R}^{C\times C}$ is computed through matrix multiplication. The output of the CTMB $F_{\text{out}}$ represents the color-corrected feature, and the bright light frame with corrected color tone can be derived from $F_{\text{out}}$ and the exposure parameter from MCM. This allows CTCM to generate global color correction via channel-wise self-attention while enhancing local color adjustments through convolutional operations.
3.4 Triplet Supervision for Radiance Field in Dark
There are three major challenges in supervision under low light conditions: First, events are too noisy to provide accurate supervision alone. Second, pseudo-bright frames generated by CTCM cannot provide robust details due to the data gap between pre-trained knowledge and actual training scenarios. Third, in low light environments, captured dark frames are easier to motion blur as a longer exposure time is required. Thus, a sharpening module is needed because the pseudo-bright frames may also contain motion blur when given blurred dark frames.
@@ 3.4 Triplet Supervision for Radiance Field in Dark
At a time window with range $t_{1}$ and $t_{2}$, given two low-light dark frames $I^{t_{1}}_{low}$ and $I^{t_{2}}_{low}$ captured at these two timestamps, and the ground truth accumulated events $E_{gt}$ between the time window, the Color Tone Consistency Module, CTCM, generates two estimated bright frame $B^{t_{1}}$, and $B^{t_{2}}$, where $(B^{t_{1}},B^{t_{2}})=\text{CTCM}(I^{t_{1}}_{low},I^{t_{2}}_{low},E_{gt})$. We then use the rendered frame from the GS model to compute $\mathcal{L}2$ loss with an estimated bright frame $B$. Therefore, the frame-based holistic loss is defined as:
@@
Event-level Granular Supervision. Granular details are ignored by the frame camera when captured in the dark due to limited dynamic range. Furthermore, the estimated pseudo-bright frame from CTCM can only give a holistic view, ignoring the same region of sharp and accurate details. Event streams captured by the event camera play a critical role in supervising details and minors because they are high in temporal resolution, so details between frames can be captured, and they are wide in dynamic range, so unrevealed information can be seen. We use a y-noise filter [7] $Y$ to form ground truth event $E_{gt}=Y(E_{raw})$ from raw captured events $E_{raw}$. To learn $E_{pred}$ with the formulated clean event signals $E_{gt}$, we adopt the classic event supervision loss from [21], where $N$ is the number of viewpoints:
@@
Mixed-Modality Sharpening Supervision. Because the generated pseudo-bright frames from CTCM may contain motion blur, applying a strict $\mathcal{L}2$ loss leads to the rendered results being blurry as well. To correct motion blur in the rendering process, we propose a mixed-modality sharpening loss. The key idea is: though individual generated bright frames from CTCM contain motion blur, consecutive frames retain temporal consistency - motion blur and sharp areas of nearby frames below to similar area and thus can be compensated by subtraction (Fig. 4 green boxes). Mathematically speaking, the subtraction (Fig. 4 absolute difference) between two generated bright frames is sharp and stands for the pixel shift in spatial. Thus, it can be a good guide for the spatial shift to rendered GS results. Therefore, we map the sharp subtraction results using the logarithmic function $L$, defined in section 3.2, to formulate the computed event signal and then supervise the predicted events from the 3D GS network. This effectively eliminates the blurring effect and produces sharper results during rendering. The mixed-modality sharpening loss $\mathcal{L}_{mix}$ is defined as:
$\begin{split}\mathcal{L}_{mix}(I^{t_{1}},I^{t_{2}},&B^{t_{1}},B^{t_{2}})=\\ \frac{1}{N}\sum_{i=1}^{N}(L(I^{t_{1}}-I^{t_{2}})&-L(B^{t_{1}}-B^{t_{2}}))^{2}.\end{split}$ (8)
@@
Noise filtering using the $y$-noise filter effectively removes interfering signals, especially in dark and low-light, while the triplet supervision keeps the training stable across varying conditions. This stability ensures consistent rendering performance, regardless of data source, and is supported by ablation studies in the supplementary materials.
4 Experiments

### v2

@@ Abstract
In low-light environments, conventional cameras often struggle to capture clear multi-view images of objects due to dynamic range limitations and motion blur caused by long exposure. Event cameras, with their high-dynamic range and high-speed properties, have the potential to mitigate these issues. Additionally, 3D Gaussian Splatting (GS) enables radiance field reconstruction, facilitating bright frame synthesis from multiple viewpoints in low-light conditions. However, naively using an event-assisted 3D GS approach still faced challenges because, in low lights, events are noisy, frames lack quality, and the color tone may be inconsistent. To address these issues, we propose Dark-EvGS, the first event-assisted 3D GS framework that enables the reconstruction of bright frames from arbitrary viewpoints along the camera trajectory. Triplet-level supervision is proposed to gain holistic knowledge, granular details, and sharp scene rendering. The color tone matching block is proposed to guarantee the color consistency of the rendered frames. Furthermore, we introduce the first real-captured dataset for the event-guided bright frame synthesis task via 3D GS-based radiance field reconstruction. Experiments demonstrate that our method achieves better results than existing methods, conquering radiance field reconstruction under challenging low-light conditions. The code and sample data are included in the supplementary material.
I Introduction
Low-light radiance field reconstruction plays a crucial role in various real-world applications, including nighttime photography [25, 53, 28], surveillance [24], and autonomous driving [14, 39, 27]. However, traditional frame-based cameras struggle to capture high-quality images in dark environments due to their limited dynamic range and reliance on long exposure times. These constraints result in significant noise and motion blur, making it difficult to capture clear, bright frames of objects from multiple viewpoints in low-light conditions.
Event cameras are a novel type of sensor that captures brightness changes asynchronously per pixel, marking a shift from conventional frame-based imaging [44, 5, 1, 13, 26]. With high dynamic range (HDR) and high temporal resolution capabilities, they can simultaneously capture bright and dark regions with no motion blur, enabling the recovery of visual details often lost in frame-based imaging, and even supporting radiance field reconstruction, particularly under low-light conditions. Event-based video reconstruction methods are proposed [33, 9, 23] to output consecutive frames from events. Nevertheless, either they are limited to grayscale images [33, 9], or primarily focus on the short exposure problem [23]. EvLight [22] leverages event data for single-image enhancement rather than focusing on video, and may exhibit partial chromatic aberrations. Event data have also been applied to radiance field reconstruction, but existing approaches [40, 34, 42] are only designed for normal light conditions (around 300 lux).
@@ I Introduction
Reconstructing a radiance field in the dark is particularly challenging due to several factors. First, the captured signals lack quality: dark frames contain extremely low intensity and brightness, while event streams become significantly noisier and sparser because of photon scarcity (Fig. 1 (a) and (c)). This makes both frame-only methods (e.g., 3D GS [20]) and event-based radiance field approaches [40, 34, 42] ineffective under such conditions. Second, ensuring color tone consistency across rendered views is non-trivial. Existing event-based video reconstruction works [33, 9, 23] either output grayscale sequences or focus on short-exposure enhancement, and thus cannot be directly adapted for radiance field reconstruction in dark environments. Third, accurate camera pose estimation and data availability present additional obstacles. On the one hand, structure-from-motion pipelines struggle in dark environments because frames lack sufficient intensity for reliable feature extraction, and no mature event-based COLMAP alternative exists that can robustly handle noisy, sparse events. On the other hand, the absence of dedicated datasets further limits progress: most prior studies rely on synthetic short/long exposure pairs and simulated events [18], which fail to capture the noise characteristics, dynamic lighting, and inconsistencies of real-world low-light data. As a result, even when accurate ground-truth poses are provided in controlled experiments, such settings do not reflect real-world deployments, and models trained on synthetic data generalize poorly to real-world low light datasets and environments.
In this paper, we propose Dark-EvGS by integrating the event camera and 3D Gaussian Splatting (GS) [20], a radiance field reconstruction technique, enabling bright frame reconstruction from arbitrary viewpoints along the camera trajectory. The high-dynamic sensing capability of event cameras enables Dark-EvGS to have a precise perception of camera motion, which can provide valuable guidance for radiance field reconstruction performance (Fig. 1 (b)) and render bright frames from multiple viewpoints of objects in low-light conditions (Fig. 1 (d)).
@@
  • We introduce Dark-EvGS, the first event-guided bright frame synthesis method via 3D GS-based radiance field reconstruction from arbitrary viewpoints in the dark.
@@
  • We collect the first real-world event-based low-light radiance field reconstruction dataset with paired low-light, bright-light frames, event streams, and corresponding camera poses. We will release the full dataset and code to facilitate future research.
II Related Work
@@
Most NeRF and 3D GS-based methods use frame data from traditional cameras [29, 20]. Few works have attempted to reconstruct radiance fields in dark and low-light environments due to the limited dynamic range of a frame-based camera. Zhang et al.[51] attempt to solve robot exploration in the dark using 3D Gaussians Relighting, but a constant illumination placed in front of the robot is necessary. Mildenhall et al.[29] use NeRF to reconstruct the radiance field in the dark by treating it as a denoising problem from HDR frames. However, it requires an HDR sensor in real applications and does not explore the effect of motion blur in low-light conditions. Some works [19, 41, 40, 11] have devoted to event-based radiance field reconstruction. But they all reconstruct under normal lighting conditions. Thus, limited work has been devoted to leveraging the dynamic range and temporal resolution properties of an event camera to solve such an issue.
Event-Based Video Reconstruction. Many works have been devoted to the field of events to video reconstruction [33, 32, 9]. However, these methods do not work well for reconstructing frames from low-light events. EvLowLight [23] attempts to reconstruct bright-light video by incorporating event data. However, the focus is primarily on the short-to-long exposure problem rather than the dark-to-bright reconstruction. In their case, the darkness of the frames is caused by short exposure times, not actual low-light conditions, meaning the scene remains well-lit from the event camera’s perspective. EvLight [22] utilizes event data to enhance single images rather than addressing video reconstruction, and its results may still exhibit partial chromatic aberrations. In contrast, our task focuses on using an event camera to reconstruct bright radiance fields in real low-light environments. Others [50, 31] have tried to use an event to guide video reconstruction in the dark. However, estimating absolute intensity values in a video solely from brightness changes recorded in events is a highly ill-posed problem [23]. Thus, directly applying any of these models to our task is not feasible.
@@
III Methodology
III-A Preliminary on 3D Gaussian Splatting
Figure 2: Overview of the Dark-EvGS pipeline for radiance field reconstruction in the dark. We obtain dark frames, events, and camera parameters using an event camera under low light. The frame encoder and event encoder will extract features, which will then be forwarded to multimodal coherence modeling. The proposed Color Tone Consistency Module takes the features and decodes them into pseudo bright frames with consistent color tone (described in Sec. III-C). We used pre-trained weights from [23] for the initialization of encoders and the multimodal coherence modeling block. The ice icon denotes frozen weights, and the fire icon denotes optimizing weights. Via the 3D GS model, we render bright and shape frames when given the corresponding camera position and pose, and formulate (blue arrow) the rendered results to predicted event signals. The same process applies when formulating captured events to ground truth supervision signals. To supervise accurate training, the proposed triplet-level loss (red lines) provides a holistic view while keeping the rendered results sharp and accurate in detail. As a result, our method enables high-quality radiance field reconstruction in the dark.
@@ III-A Preliminary on 3D Gaussian Splatting
III-B Noisy Events in Dark as Supervisory Signals
@@
Events triggered in low-light environments are much noisier [18, 10]. To address noise, which is often more prevalent in microsecond-level event data [43], we preprocess the event stream using the $y$-noise filter from [12]. As shown in Fig. 3 (a), raw events captured in the dark contain extreme noise and randomness. After applying the noise filter, the cleaned event stream becomes more helpful as a supervision signal, as shown in Fig. 3 (b). Noise filtering is essential for accurate view rendering, as shown by the ablation study section.
Our objective, depicted in Fig. 2, is to reconstruct a radiance field using differentiable 3D Gaussian functions $G$, guided by event and corresponding blurred dark frames. We accumulate the ground truth event signal between timestamps $t_{1}$ and $t_{2}$ to form the supervisory signal using Eq. 5 by aggregating the polarities of all events occurring between times $t$ and twt-w, indexed by their pixel location.
@@
Figure 3: Frame-based cameras are unable to capture enough signals for residence reconstruction in the dark. Event cameras, on the other hand, can capture ample signals, including ignored details by frame-based cameras (a). However, these raw events captured are rather noisy and random (a), which is one of the biggest challenges for existing methods [23, 40, 34, 33]. To solve this problem, we proposed a utilization for noisy events for accurate and clean supervisory signals (b).
III-C Color Tone Consistency Module
Pure event utilization and supervision are insufficient for radiance field reconstruction on real data [15, 40], let alone reconstruction under low-light environments. Therefore, prior knowledge is required in the 3D GS training pipeline. Following EvLowLight [23], we deploy feature encoders for both frame and event modality along with the Multimodal Coherence Modeling block (MCM), with corresponding pre-trained weights for initialization. We then propose the Color Tone Consistency Module (CTCM), which consists of a proposed Color Tone Matching Block (CTMB) along with the decoders from MCM. CTMB takes the decoded frames during the decoding process and corrects the color tone. Finally, the CTCM outputs the predicted pseudo frames using the corrected frames and the exposure parameter predicted from the MCM. Specifically, the CTMB, inspired by recent advancements in transposed self-attention [37], takes the last feature map $F\in\mathbb{R}^{H\times W\times C}$, and derives query $Q$, key $K$, and value $V$ representations using a 1×1 convolution followed by a depth-wise convolution. Subsequently, $Q$ and $V$ are reshaped, and an attention map $M\in\mathbb{R}^{C\times C}$ is computed through matrix multiplication. The output of the CTMB $F_{\text{out}}$ represents the color-corrected feature, and the bright light frame with corrected color tone can be derived from $F_{\text{out}}$ and the exposure parameter from MCM. This allows CTCM to generate global color correction via channel-wise self-attention while enhancing local color adjustments through CNNs.
III-D Triplet Supervision for Radiance Field in Dark
There are three major challenges in supervision under low light conditions: First, events are too noisy to provide accurate supervision alone. Second, pseudo-bright frames generated by CTCM cannot provide robust details due to the data gap between pre-trained knowledge and actual training scenarios. Third, in low light environments, captured dark frames are easier to motion blur as a longer exposure time is required [54, 2]. Thus, a sharpening module is needed because the pseudo-bright frames may also contain motion blur when given blurred dark frames.
@@ III-D Triplet Supervision for Radiance Field in Dark
At a time window with range $t_{1}$ and $t_{2}$, given two low-light dark frames $I^{t_{1}}_{low}$ and $I^{t_{2}}_{low}$ captured at these two timestamps, and the ground truth accumulated events $E_{gt}$ between the time window, the Color Tone Consistency Module, CTCM, generates two estimated bright frames $B^{t_{1}}$, and $B^{t_{2}}$, where $(B^{t_{1}},B^{t_{2}})=\text{CTCM}(I^{t_{1}}_{low},I^{t_{2}}_{low},E_{gt})$. We then use $N$ rendered frames from the GS model to compute $\mathcal{L}2$ loss with an estimated bright frame $B$, where $N$ is the number of viewpoints. Therefore, the frame-based holistic loss is defined as:
@@
Event-level Granular Supervision. Granular details are ignored by the frame camera when captured in the dark due to limited dynamic range. Furthermore, the estimated pseudo-bright frame from CTCM can only give a holistic view, ignoring the same region of sharp and accurate details. Event streams captured by the event camera play a critical role in supervising details and minors because they are high in temporal resolution, so details between frames can be captured, and they are wide in dynamic range, so unrevealed information can be seen. We use a y-noise filter [12] $Y$ to form ground truth event $E_{gt}=Y(E_{raw})$ from raw captured events $E_{raw}$. To learn $E_{pred}$ with the formulated clean event signals $E_{gt}$, we adopt the classic event supervision loss from [34], where $N$ is the number of rendering viewpoints:
@@
Mixed-Modality Sharpening Supervision. Because the generated pseudo-bright frames from CTCM may contain motion blur, applying a strict $\mathcal{L}2$ loss leads to the rendered results being blurry as well. To correct motion blur in the rendering process, we propose a mixed-modality sharpening loss. The key idea is: though individual generated bright frames from CTCM contain motion blur, consecutive frames retain temporal consistency - motion blur and sharp areas of nearby frames belong to similar area and thus can be compensated by subtraction (Fig. 4 green boxes). Mathematically speaking, the subtraction (Fig. 4 absolute difference) between two generated bright frames is sharp and stands for the pixel shift in spatial. Thus, it can be a good guide for the spatial shift to rendered GS results. Therefore, we map the sharp subtraction results using the logarithmic function $L$, to formulate the computed event signal and then supervise the predicted events from the 3D GS network. This effectively eliminates the blurring effect and produces sharper results during rendering. Here, $I^{t_{1}}$ and $I^{t_{2}}$ denote two rendered intensity images from the 3D Gaussian Splatting model at timestamps $t_{1}$ and $t_{2}$, respectively, while $B^{t_{1}}$ and $B^{t_{2}}$ are the corresponding pseudo-bright frames generated by the CTCM. The operator $L(\cdot)$ represents the logarithmic intensity mapping used to compute log-intensity differences. The index $i$ enumerates frames, and $N$ denotes the total number of frames. The mixed-modality sharpening loss $\mathcal{L}_{mix}$ is defined as:
$\begin{split}\mathcal{L}_{mix}(I^{t_{1}},I^{t_{2}},&B^{t_{1}},B^{t_{2}})=\\ \frac{1}{N}\sum_{i=1}^{N}((L(I^{t_{1}})-L(I^{t_{2}}))&-(L(B^{t_{1}})-L(B^{t_{2}})))^{2}.\end{split}$ (8)
@@
Noise filtering using the $y$-noise filter effectively removes interfering signals, especially in dark and low-light, while the triplet supervision keeps the training stable across varying conditions. This stability ensures consistent rendering performance, regardless of data source.
IV Experiments